OpenStudy (priyar):

The reduction potential of hydrogen half cell will be negative if:

OpenStudy (priyar):

if p(H2) =a and [H+] = b then options are: a=1atm,b=1M a=2atm b=1M a=2atm b=2M a=1atm b=2M

OpenStudy (priyar):

@Rushwr @Cuanchi

OpenStudy (cuanchi):

do you have any idea? are you familiar with the Nernst equation?

OpenStudy (cuanchi):

look at SAMPLE EXERCISE 20.13 in this page http://wps.prenhall.com/wps/media/objects/3313/3392587/blb2006.html

OpenStudy (priyar):

@Cuanchi yes, i am familiar with Nernst eq. but i have a doubt what is the anode here?

OpenStudy (priyar):

@samigupta8 @imqwerty

OpenStudy (samigupta8):

@priyar it's just a half cell here so no anode and cathode is platinum electrode in which hydrogen is bubbled at 1 atm and the solution is 1M H+ concentration

OpenStudy (samigupta8):

What i m worried about is how will we get a negative potential of hydrogen through the value of b that u hve provided here ..it can be negative only if [H+] conc was lying in between 0 to 1(greater than 0 and less than 1)

OpenStudy (cuanchi):

You have to consider also the Pressure and calculate the Q. when the Q is smaller than 0 the E will be positive because the loq Q will be negative and if the Q > 0 , the log Q is positive and the E of the cell will be negative. the anode is negative and is where the oxidation occurs In this problem the anode is the standard H electrode and the cathode will be the non-standard one

OpenStudy (samigupta8):

Really @cuanchi

OpenStudy (samigupta8):

Won't it be like the way i suggested

OpenStudy (priyar):

@ganeshie8 @photon336

OpenStudy (priyar):

@ParthKohli

OpenStudy (priyar):

@Cuanchi please help

OpenStudy (priyar):

The reaction will be: H+ + e- --> 1/2 H2 Eo = Eright - Eleft = x -0 =x and Q= p(H2)^1/2/[H+] what will be Ecell (that is the RHS of Nernst equation)?

OpenStudy (priyar):

@Vincent-Lyon.Fr

OpenStudy (priyar):

@Kainui

OpenStudy (cuanchi):

a=2atm b=1M

OpenStudy (samigupta8):

I agree with @cuanchi...

OpenStudy (priyar):

@Cuanchi can you explain how you solved?

OpenStudy (cuanchi):

\[Q= \frac{ ([H+] electrode1)^{2} \times P _{H2} electrode2 }{([H+] electrode2)^{2} \times P _{H2} electrode1 }= \frac{ 1\times P _{H2} electrode2}{ [H+] electrode2)^{2} \times 1 }\] we agree that the electrode 1 is the standard electrode \[Q=\frac{ P _{H2} }{ [H+]_{2}^{2} }\] when the Q is smaller than 0 the E will be positive because the loq Q will be negative and if the Q > 0 , the log Q is positive and the E of the cell will be negative.

OpenStudy (priyar):

what about Eo of cell (standard potential ) ? Nernst Eq has both Ecell and Eo cell.. can you tell me what will the value of both in this eq? coz thats where my doubt lies..

OpenStudy (priyar):

@Cuanchi

OpenStudy (cuanchi):

Eo is equal to "0" if the two half cells are the same composition (in standard conditions the concentration of the cells has to be 1M and the pressure 1atm by definition of standard conditions). When you have a concentration cells, that means the only difference between the two cells is the concentration of the solutions or the pressure of the gases, the standard potential "Eo = 0" then you need the Nernst equation to calculate the E that it is dependent of the Q (Q=1 => log Q =0). When two different half cells are in standard condition (1M, 1atm) the E = Eo and it is independent of the Q . If you have both, different half cells composition and they are not at standard conditions you use the Nernst eq and you calculate the Eo with the standard potential of the half cells and the Q with the concentrations of products and reactants.

OpenStudy (priyar):

so these are concentration cells? so we use nernst eq with Eo=0 right? Thanks!

OpenStudy (cuanchi):

yes!!! definitely they are!